# Learning Simulatable Models of Cloth with Spatially-varying Constitutive Properties

### v2

@@ 5 Conclusion
We presented Mass-Spring Net, a surrogate model trained with a novel force-impulse loss and training curriculum that can learn complex, spatially-varying constitutive properties of a piece of simulated cloth without relying on a reference mesh. We demonstrated that our method can reproduce the observed behavior of both homogeneous and heterogeneous cloth sheets better than previous neural network approaches. It generalizes more readily to new scenarios, including those with previously unobserved contacts. Further, we showed that Mass-Spring Net’s mass-spring-based architecture is immune to membrane locking that plagues FEM-based simulations.
@@
Our method has some limitations. While mass-spring lattices can represent a wide-range of material properties, it is uncertain if they provide universal approximating power for constitutive behavior. In our setup we used specific spring network topologies, with some examples not using bending springs connecting to the 2-ring around a vertex. This may explain some of the minor mismatches we see in the dynamic and static behavior. The topology itself could be learned for better performance and better resolution of the spatial variation. Finally, to learn spatially varying material properties, all parts of the cloth should be “‘sufficiently excited” in the training data. Better methods to achieve this than dropping the cloth in various ways could be designed, to improve the recovery of material properties.
@@
Despite these limitations, we were pleasantly surprised by how well mass-spring systems work when trained with FEM simulation data, and their surprising ability to avoid membrane locking artifacts even when they are present in some of the training examples. Thus, when trained by our offline process, Mass-Spring Net could provide effective bridge between sophisticated and slow FEM simulators and the needs of real-time applications in games and VR.
@@ Appendix A Synthetic Motion Clip Set Construction
Other than particle positions and velocities over time, for each rollout we register 1. the particles fixed in space which is the boundary condition discussed in Sec. 4.1; 2. the mass of the cloth. For a simulated, dynamic source cloth this can be accessed by querying the area density and surface area of the cloth from the simulator used to generate data; for real-world objects we expect users to have the means of measuring the object’s weight and from there they can obtain the mass of the cloth. 3. physical dimension of the cloth, i.e. length and width. We assume the cloth always take a rectangular shape.
@@
Construction point cloud clip dataset. Following the approach adopted by prior work on system identification with differentiable simulators [Ma et al., 2023, Gärtner et al., 2022] in training, we partition each rollout into non-overlapping “clips” with each clip containing time steps of simulation data, and make the surrogate system track the source system’s motion throughout a clip on each training iteration. For training mass-spring net, we set , so each rollout gives clips. With rollouts we get clips, or time steps of kinematic data. To reduce memory footprint and disk space usage, we downsize the dataset by sampling of all the clips generated, and that leaves us time steps of training data. We then take clips from the last of each rollout and subsample of the clipsto yield time steps of low-velocity training motion data.
@@ Appendix B Implementation Details and Hyperparameters
B.1 Implementation Details
As in Ma et al. [2023] we implement our neural constitutive model in PyTorch [Paszke et al., 2019], and we implement contact resolution and forward dynamics in WARP [Macklin, 2022b]. Both forward and backward propagation through our surrogate model are fast, since CUDA tensors from PyTorch can be directly used by WARP, and we have exploited WARP’s parallelism by allowing simulation of multiple clips to happen in parallel. All experiments are run on NVIDIA V100’s with GB of GPU memory.
@@ B.2 Hyperparameters
Impulse loss vs. force loss. When training Mass-Spring Net on kinematic data sampled from mass-spring based cloth, we used a combination of force loss and impulse loss; specifically, we set and in both the stiffness and the damping pass. On the other hand, for training on data sampled from mesh-based cloths, in the stiffness pass we used the impulse loss exclusively with , and in the damping pass we used the force loss exclusively with . We believe the optimal combination of the force and impulse term depends on (i) the source of training data and (ii) whether we are estimating springs’ elastic stiffness or damping properties. What is particularly interesting is that for FEM-mesh-based training data, using impulse loss is crucial for learning stiffness, yet both impulse and force loss are crucial for obtaining good damping estimates. We hypothesize that impulse loss which tracks error in force accumulation over time is more resistant to noises induced by taking the second derivative with respect to particle positions, therefore more suitable for learning from low-velocity data in the stiffness pass. On the other hand, force loss is sensitive to minor perturbations induced by taking derivatives. While this is the case for learning from meshes that bear a different resolution from the surrogate, we see that in the case of learning from mass-spring sheet with the same resolution as the surrogate, both impulse and force loss enable effective learning in either the stiffness or the damping pass.
@@
Hyperparameter choices for the reconstruction accuracy experiment. We train the mass-spring net for iterations, including iterations in the first (stiffness) pass and in the second (damping) pass. We treat each -step-long motion clip as a sample, and we simulate with batch size of clips in forward and backpropagation. Given that the training set contains clips, training for iterations is equivalent to training for epochs. For the loss function, we use , . We notice that including the force loss is critical for narrowing the distribution of stiffness and damping estimates near ground-truth values.
@@
Hyperparameter choices for the membrane locking experiment. To model actual bending and show that our bending springs can model the proper bending in a source FEM mesh, we included bending springs in the surrogate. Also, we included both diagonal springs to eliminate the possibility of introducing pseudo-stiffness in the misaligned case due to asymmetry. We trained MGN [Pfaff et al., 2020] for steps over 36 hours. The number of training iterations we used is much lower than the iterations Pfaff et al. [2020] used to train MGN on the flag_simple example reported in their paper. We trained for fewer steps because (i) the low-resolution mesh is more coarse than the flag_simple example, as our mesh contains roughly only vertices as the mesh they used; (ii) time constraints, as training for steps would require more than days on an NVIDIA RTX3090 GPU. The excessively long training time can be partially attributed to lacking of parallelism in the training curriculum, e.g. simulations can only be run in batches of rollout; but also can be attributed to not leveraging sufficient physical priors for inferring constitutive properties, similar to what GNN suffers from (which we will discuss below). Those being said, for a qualitative assessment to demonstrate that MGN can be vulnerable to membrane locking in training data, the results from training after iterations should be sufficient, because we expect that after training for more iterations the results will be more visually close to the low-resolution source system’s equilibrium configuration, and we do not see evidence of MGN learning to resist membrane locking throughout the course of training.
@@
Hyperparameter choices for the generalization experiment in which we subject Mass-Spring Net and the baselines under novel initial and boundary conditions. The Neural ODE baseline contains hidden layers, with neurons in each hidden layer, and we use the exponential ReLU activation after each hidden layer. We train Mass-Spring Net for iterations in the stiffness pass, and for iterations in the damping pass, until loss converges. we train Neural ODE for iterations, with the same loss function as used for Mass-Spring Net, on a single pass, and we have made sure that training loss had converged. We train GNN for iterations. But note that training of GNN works a bit differently: (i) we train on rollouts instead of rollouts, but guarantee that the amount of data, measured in terms of number of time steps is the same as using rollouts; (ii) instead of using -step long clips we use -step long clips. Each clip is x-regularly downsampled from steps of simulation data throughout seconds of simulation. (iii) training uses batch size of clip. We have tried training GNN on rollouts partitioned into -step-long clips, but GNN’s performance of simulating long test rollouts became worse. We believe using -step-long clips didn’t work because these clips consist of stepwise data that are contiguous in time, meaning they condition GNN to reconstruct short rather than long rollouts well. On the other hand, using more rollouts ( vs ) did not help, because GNN may had difficulties with capturing the underlying physics, i.e. neural constitutive properties and including more training data only complicated the objective landscape which GNN already struggled to learn. It is also interesting to note that training GNN takes an excessively large amount of time: hours, compared to roughly hours for our method and Neural ODE. We attribute this to the fact that GNN is tailed to learning generic granular materials as opposed to a piece of cloth with a fixed topological structure as known prior. Specifically, GNN requires users to set up a hyperparameter named “radius”, which defines a spherical region over which message-passing between particles take place. Granular materials like sands generally only requires a small radius for message passing, because each particle only physically interacts with particles that are spatially close to it. This however is not the case for a piece of cloth: consider a scenario in which two vertices lie on two opposite sides of the cloth. if the cloth is folded over such that the two sides become close to each other, until the two sides touch each other there should be minimal to no interaction between the two particles, because they are topologically far apart. Yet without the topology of the cloth as a known prior, GNN would make message passing happen between the two particles simply based on the fact that they are spatially close. Such redundant message passing introduces significant performance overheads. Turning to the loss function: we use the impulse loss exclusively, and we found that force loss is not particularly useful in this case. We believe this is because when the source cloth is constructed from triangular meshes, elastic and damping forces that act on each particle comes from edges which has no one-to-one correspondence to springs in the surrogate. Therefore learning per-spring elastic and damping force from the net force applied onto each particle becomes difficult. Impulse loss on the other hand is more robust to the change in geometric construction.
@@ Appendix C Additional Results
C.1 Generalizability to Novel Initial and Boundary Conditions
@@
Fig. 12 shows the architecture of the MLP-based Neural ODE that we compare with Mass-Spring Net as a baseline. Without topological structure of the sheet explicitly injected into the architecture as a prior knowledge, the neural network needs to learn it by itself.
@@
Fig. 13 shows the source and surrogate sheets reaching four different steady states, after four -seconds-long rollouts started from four different initial and boundary conditions. Note that for each rollout, all surrogates start from the same condition as the source. Notice that our method yields significantly lower steady state errors than the baselines. Neural ODE’s prediction completely disintegrates after seconds; GNN does better but still, deviation from the ground-truth is significant. Nevertheless, our method still has room for improvement. Notice that in the first row in Fig. 13, the bottom-left corner of the cloth is poorly fitted. Also notice that in the third row the mass-spring based cloth is further stretched downward than the ground-truth.
More analysis on GNN’s performance of reconstructing long rollouts. Compared with GNN, our method is able to maintains stable reconstruction for extended () rollouts, demonstrating greater robustness. So why does GNN perform comparably to Mass-Spring Net for short () rollouts but their performance diverges for longer rollouts? We believe difference stems from our method’s focus on cloth simulation, while GNN targets generic granular materials. Without leveraging the time-invariant topological structure of a cloth, GNN attempts to model dense particle-wise interactions at each time step, thereby complicating the learning process and prolonging training time. Note that GNN’s performance on -long rollouts is comparable to the number reported in NCLaw [Ma et al., 2023] which also evaluated GNN on short rollouts.
C.2 Ablation Studies
@@
Dual-pass vs single-pass training. An obvious alternative to the dual-pass training curriculum would be to learn stiffness and damping constants at once, in a single pass. We conducted a single-pass train on mass-spring cloth data and found that compared with results shown in Fig. 7 while the RMSE of k estimates only increased slightly from to , the RMSE of b estimates increased significantly from to . This is expected because of the entanglement between elastic and damping force as discussed in Sec. 3.3.
@@
Force-and-impulse loss vs position loss. Position-based loss which directly penalizes step-wise Euclidean distance between target and predicted particle positions is commonly used in neural constitutive modelling [Bianchi et al., 2003, Ma et al., 2023, Zong et al., 2023]. We tried learning stiffness with position loss on mass-spring cloth data. We use regularization strength , since position loss tends to be larger than force or impulse loss. Other hyperparameters are the same as when using force-and-impulse loss. We notice that with position loss, the RMSE of stiffness estimates increased significantly from to , and reconstruction RMSE of test motions increased from to . We believe what makes force-and-impulse loss or specifically, force loss working really well when the source cloth is made up of masses and springs is that it is highly sensitive to stepwise difference between predicted and target elastic and damping forces. Position loss like impulse loss effectively integrates errors over time, and the accumulation of per-step errors can make the neural model find it difficult to estimate system parameters on a finer scale.

### v3

@@ 5. Conclusion
We presented SpringTime, a surrogate model trained with a novel force-impulse loss and training curriculum that can learn complex, spatially-varying constitutive properties of a piece of simulated cloth without relying on a reference mesh. We demonstrated that our method can reproduce the observed behavior of both homogeneous and heterogeneous cloth sheets better than previous neural network approaches. It generalizes more readily to new scenarios, including those with previously unobserved contacts. Further, we showed that SpringTime’s mass-spring-based architecture is fairly robust against membrane locking which plagues FEM-based simulations.
@@
Our method has some limitations. While mass-spring lattices can represent a wide-range of material properties, it is uncertain if they provide universal approximating power for constitutive behavior. The topology itself could be learned for better performance and better resolution of the spatial variation. Finally, to learn spatially varying material properties, all parts of the cloth should be “‘sufficiently excited” in the training data. Better methods to achieve this than dropping the cloth in various ways could be designed, to improve the recovery of material properties.
@@
Despite these limitations, we were pleasantly surprised by how well mass-spring systems work when trained with FEM simulation data, and their surprising ability to avoid membrane locking artifacts even when they are present in some of the training examples. Thus, when trained by our offline process, SpringTime could provide effective bridge between sophisticated and slow FEM simulators and the needs of real-time applications in games and VR.
@@ Appendix A Synthetic Motion Clip Set Construction
Other than particle positions and velocities over time, for each rollout we register (1) the particles fixed in space which is the boundary condition discussed in Sec. 4.1; (2) the mass of the cloth. For a simulated, dynamic source cloth this can be accessed by querying the area density and surface area of the cloth from the simulator used to generate data; for real-world objects we expect users to have the means of measuring the object’s weight and from there they can obtain the mass of the cloth. (3) physical dimension of the cloth, i.e. length and width. We assume the cloth always take a rectangular shape.
@@
Construction of the point cloud clip dataset. Following the approach adopted by prior work on system identification with differentiable simulators (Ma et al., 2023; Gärtner et al., 2022) in training, we partition each rollout into non-overlapping “clips” with each clip containing steps of simulation data, and make the surrogate system track the source system’s motion throughout a clip on each training iteration. For training mass-spring net, we set , so each rollout gives clips, where of rollout gives steps of data at timestep size of . So from rollouts we get clips, or time steps of kinematic data. To reduce memory footprint and disk space usage, we regularly downsample time steps– taking in every steps, and that leaves us time steps of training data. We then take clips from the last of each rollout to yield time steps of low-velocity training motion data.
@@ Appendix B Implementation Details and Hyperparameters
B.1. Implementation Details
As in Ma et al. (2023) we implement our neural constitutive model in PyTorch (Paszke et al., 2019), and we implement contact resolution and forward dynamics in WARP (Macklin, 2022a). Both forward and backward propagation through our surrogate model are fast, since CUDA tensors from PyTorch can be directly used by WARP, and we have exploited WARP’s parallelism by allowing simulation of multiple clips to happen in parallel. All experiments are run on NVIDIA V100’s with GB of GPU memory.
@@ B.2. Hyperparameters
Impulse loss vs. force loss. When training SpringTime on kinematic data sampled from mass-spring based cloth, we used a combination of force loss and impulse loss; specifically, we set and in both the stiffness and the damping pass. On the other hand, for training on data sampled from mesh-based cloths, in the stiffness pass we used the impulse loss exclusively with , and in the damping pass we used the force loss exclusively with . We believe the optimal combination of the force and impulse term depends on (i) the source of training data and (ii) whether we are estimating springs’ elastic stiffness or damping properties. What is particularly interesting is that for FEM-mesh-based training data, using impulse loss is crucial for learning stiffness, yet both impulse and force loss are crucial for obtaining good damping estimates. We hypothesize that impulse loss which tracks error in force accumulation over time is more resistant to noises induced by taking the second derivative with respect to particle positions, therefore more suitable for learning from low-velocity data in the stiffness pass. On the other hand, force loss is sensitive to minor perturbations induced by taking derivatives. While this is the case for learning from meshes that bear a different resolution from the surrogate, we see that in the case of learning from mass-spring sheet with the same resolution as the surrogate, both impulse and force loss enable effective learning in either the stiffness or the damping pass.
@@
Hyperparameter choices for the reconstruction accuracy experiment. We train SpringTime for iterations, including iterations in the first (stiffness) pass and in the second (damping) pass. We treat each -step-long motion clip as a sample, and we simulate with batch size of clips in forward and backpropagation. Given that the training set contains clips, training for iterations in the stiffness pass is equivalent to training for epochs. For the loss function, we use , . We notice that including the force loss is critical for narrowing the distribution of stiffness and damping estimates near ground-truth values.
@@
Hyperparameter choices for the membrane locking experiment. To model actual bending and show that our bending springs can model the proper bending in a source FEM mesh, we included bending springs in the surrogate. Also, we included both diagonal springs to eliminate the possibility of introducing pseudo-stiffness in the misaligned case due to asymmetry. We trained MGN (Pfaff et al., 2020) for steps over 48 hours. The number of training iterations we used is much lower than the iterations Pfaff et al. (2020) used to train MGN on the flag_simple example reported in their paper. We trained for fewer steps because (i) the low-resolution mesh is more coarse than the flag_simple example, as our mesh contains roughly only vertices as the mesh they used; (ii) time is the constraining factor– training for steps would require more than hours on a server with NVIDIA V100 which is prohibitively long for our purpose. The excessively long training time can be partially attributed to lacking of parallelism in the training curriculum, e.g. simulations can only be run in batches of rollout; but also can be attributed to not leveraging sufficient physical priors for inferring constitutive properties. Those being said, for a qualitative assessment to demonstrate that MGN can be vulnerable to membrane locking in training data, the results from training after iterations should be sufficient, because we expect that after training for more iterations the results will be more visually close to the low-resolution source system’s equilibrium configuration, and we do not see evidence of MGN learning to resist membrane locking throughout the course of training.
@@
Hyperparameter choices for the generalization experiment in which we subject SpringTime and the baselines under novel initial and boundary conditions. The Neural ODE baseline contains hidden layers, with neurons in each hidden layer, and we use the exponential ReLU activation after each hidden layer. We train SpringTime for iterations in the stiffness pass, and for iterations in the damping pass, until loss converges. we train Neural ODE for iterations, with the same loss function as used for SpringTime, on a single pass, and we have made sure that training loss had converged. We train MGN for million iterations. But note that training of MGN works a bit differently: (i) we train on rollouts with -step long clips. The downsampling strategy introduced in Sec. A is applied here as well, so we get frames of data per clip. This is on the same scale as frames of data per clip used in training the flag_dynamic example in MGN Pfaff et al. (2020). (ii) training uses batch size of clip. Turning to the loss function: we use the impulse loss exclusively, and we found that force loss is not particularly useful in this case. We believe this is because when the source cloth is constructed from triangular meshes, elastic and damping forces that act on each particle comes from edges which has no one-to-one correspondence to springs in the surrogate. Therefore learning per-spring elastic and damping force from the net force applied onto each particle becomes difficult. Impulse loss on the other hand is more robust to the change in geometric construction.
@@ Appendix C Additional Results
C.1. Generalizability to Novel Initial and Boundary Conditions
@@
Fig. 12 shows the architecture of the MLP-based Neural ODE that we compare with SpringTime as a baseline. Without topological structure of the sheet explicitly injected into the architecture as a prior knowledge, the neural network needs to learn it by itself.
@@
Fig. 13 shows the source and surrogate sheets reaching four different steady states, after four -seconds-long rollouts started from four different initial and boundary conditions. Note that for each rollout, all surrogates start from the same condition as the source. Notice that our method yields significantly lower steady state errors than Neural ODE in particular, for which the prediction completely disintegrates after seconds; MGN performs comparably to SpringTime but requires considerably training time to obtain the shown results. Nevertheless, our method still has room for improvement. Notice that in the first row in Fig. 13, the bottom-left corner of the cloth is poorly fitted. Also notice that in the third row the mass-spring based cloth is further stretched downward than the ground-truth.
C.2. Ablation Studies
@@
Dual-pass vs single-pass training. An obvious alternative to the dual-pass training curriculum would be to learn stiffness and damping constants at once, in a single pass. We conducted a single-pass training session of iterations, with batch size of clips on mass-spring cloth data, and we found that compared with results shown in Fig. 7 the RMSE of stiffness constant estimates increased from to , and the RMSE of damping constant estimates increased from to . This is expected because of the entanglement between elastic and damping force as discussed in Sec. 3.3.
@@
Force-and-impulse loss vs position loss. Position-based loss which directly penalizes step-wise Euclidean distance between target and predicted particle positions is commonly used in neural constitutive modelling (Bianchi et al., 2003; Ma et al., 2023; Zong et al., 2023). We tried learning stiffness with position loss on mass-spring cloth data. We use regularization strength , since position loss tends to be larger than force or impulse loss. Other hyperparameters are the same as when using force-and-impulse loss. We notice that with position loss, the RMSE of stiffness estimates increased significantly from to . We believe what makes force-and-impulse loss or specifically, force loss working really well when the source cloth is made up of masses and springs is that it is highly sensitive to stepwise difference between predicted and target elastic and damping forces. Position loss like impulse loss effectively integrates errors over time, and the accumulation of per-step errors can make the neural model find it difficult to estimate system parameters on a finer scale.